Solve for X
x^2 + 4x + 4 = 7
(factor the left hand side)
4x + 4 = 8x
x^2 + 8x = 7
I have 2 X's and this confuses me since I can not factor them to come up with it to = 7, can someone please explain, thank you

Answer by checkley71(8403)

X^2+4X+4=7
X^2+4X+4-7=0
X^2+4X-3=0 THIS QUADRATIC DOES NOT FACTOR NORMALLY THUS WE NEED TO USE THE QUADRATIC FORMULA TO FIND THE VALUES OF X THUS
X=(-B+-SQRT[B^2-4AC])/2A
X=(-4+-SQRT[4*4-4*1*-3])/2*1
X=(-4+-SQRT[16+12])/2
X=(-4SQRT[28])/2
X=(-4+-5.29)/2
X=(-4+5.29)/2
X=1.29/2
X=.645 SOLUTION
X=(-4-5.29)/2
X=-9.29/2
X=-4.645 SOLUTION